OpenStudy (anonymous):

which three measurements could be lengths of the sides of a triangle? select one a 4cm 6cm 3cm b 4cm, 4cm, 9cm, c 4cm. 4cm, 8cm, d 4cm, 6cm, 1cm,

OpenStudy (lgbasallote):

the length of two sides must be greater than the third side for it to be a triangle. for example.. 2,2,3: it can form a triangle because 2+ 2> 3 2 + 3 > 2 however, 1,2, 3 cannot be a triangle because 1 + 2 is not greater than 3

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

yeah but im stuck between a and d

OpenStudy (lgbasallote):

in D.. D. 4, 6, 1 4 + 6 = 10 <--greater than 1 \(\checkmark\) 6 + 1 = 7 <---greater than 4 \(\checkmark\) 4 + 1 = 5 <--NOT greater than 6 X

OpenStudy (lgbasallote):

in A. A. 4,6,3 4 + 6 = 10 <--greater than 3 \(\checkmark\) 4 + 3 = 7 <--greater than 5 \(\checkmark\) 6 _ 3 = 9<--greater than 4 \(\checkmark\)

OpenStudy (lgbasallote):

the middle one is supposed to say greater than 6 not 5

OpenStudy (lgbasallote):

got it @infamous21 ?

OpenStudy (anonymous):

yeah so the answer is a @lgbasallote

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

thank you

OpenStudy (lgbasallote):

welcome